# Autodeleveraging: Impossibilities and Optimization

### v1

@@ 2.4 Autodeleveraging
    For $i\in[n]$, define $A_{i}=(B-\sum_{j=1}^{i}e_{T}(\mathfrak{p}_{\sigma(j)})_{+})_{+}$, where $e_{T}(\mathfrak{p}_{\sigma(j)})$ is the equity of the $j$th largest position under the ranking $\sigma$
@@
    Define $h\in[0,1]^{n}$ as
    $h_{\pi_{PL},\sigma(i)}=\begin{cases}1&\text{if }A_{i-1}-A_{i}=e_{T}(\mathfrak{p}_{\sigma(i)})_{+},\;A_{i}>0\\ \frac{A_{i-1}}{e_{T}(\mathfrak{p}_{\sigma}(i))}&\text{if }A_{i}=0,A_{i-1}>0\\ 0&\text{otherwise}\end{cases}$ (19)
Note this definition of $h_{\pi_{PL}}$ ensures that $\sum_{i=1}^{n}h_{\pi_{PL},i}e_{T}(\mathfrak{p}_{i})=B$, satisfying (17). On the other hand, since we only subtract positive equity positions (e.g. $j$ such that $e_{T}(\mathfrak{p}_{\sigma(j)})_{+}>0$), we satisfy (18) by construction.
@@
This haircut strategy is even simpler than the PNL-Leverage ranking and has some benefits in terms of fairness, as we will show in §6. Given a valid feasible shortfall $\theta_{\pi}D_{T}(\mathcal{P}_{n}\leq W_{T}(\mathcal{P}_{n}))$, the pro-rata haircut policy is defined as:
$h_{\pi_{PR}}(\mathfrak{p})=\frac{\theta_{\pi}D_{T}(\mathcal{P}_{n})}{W_{T}(\mathcal{P}_{n})}(e_{T}(\mathfrak{p}))_{+}$ (20)
This ranking simply says that the positive equity positions are socialized proportional to their size. Some exchanges slightly modify this formula to include leverage (e.g., Binance’s ADL Priority Index includes leverage and unrealized PNL [BinanceADL]; Aevo documentation discusses leverage-weighted priority [AevoADL]), which penalizes higher leverage positions much like the PNL-leverage ranking:
$h_{\pi_{LPR}}(\mathfrak{p})=\left(\frac{\ell e_{T}(\mathfrak{p})_{+}}{\sum_{\mathfrak{p}\in\mathcal{P}_{n}}\ell e_{T}(\mathfrak{p})_{+}}\right)h_{\pi_{PR}}(\mathcal{P}_{n})$ (21)
@@
Figure 2: ADL severity example comparing queue and pro-rata coloring. Purple shading equals the negative equity mass (deficit) while blue shading shows the haircut mass allocated to winning traders. The queue panel’s dashed blue block at rank 2 highlights residual equity when the queue method allows partial closures; exchanges that close winners fully (e.g., Hyperliquid) would shave this bar completely. Haircut mass matches deficit mass in each panel, illustrating severity $\theta=0.50$.
@@
This corresponds to an exchange guaranteeing that a user will never lose more than $\overline{h}_{i}$% of their position in a single ADL round. Moreover, an exchange might also offer an absolute guarantee to users: if your equity is positive and sufficiently large, you will never be have your equity cut to below some threshold $\underline{e}_{i}$. This can be represented by the constraint
$(1-h_{\pi,i})e_{T}(\mathfrak{p}_{i})\geq\underline{e}_{i}$ (23)
Numerous exchanges offer one or both of these guarantees on a per-ADL-round basis, including Hyperliquid, BitMEX, and Aevo [HyperliquidDocsLiquidations, BitMEXADL, AevoADL].
@@
Consider five accounts with equities $e_{T}=(10,5,1,-3,-12)$ and queue order $\sigma=(2,1,3,5,4)$ induced by the PNL-leverage scores. Let the queue policy choose severity $\theta_{\pi_{PL}}=\tfrac{1}{2}$, so the haircut budget is $B=\theta_{\pi_{PL}}D_{T}(\mathcal{P}_{n})=7.5$. The construction in (19) yields $A=(2.5,0,0,0,0)$ and haircuts $h=(0.25,1,0,0,0)$: the top winner is cut from $10$ to $7.5$, the second winner is fully wiped out, and the remaining accounts are unchanged, giving $e_{T}=(7.5,0,1,-3,-4.5)$. If, before the next ADL decision, only the losing accounts partially recover to $e_{T}=(7.5,0,1,-1.5,-2.25)$, then a second round at the same severity $\theta_{\pi_{PL}}=\tfrac{1}{2}$ autodeleverages all negative-equity accounts against the largest winner, yielding final equities $e_{T}=(3.75,0,1,0,0)$. Under the pro-rata rule (20) with the same severity $\theta_{\pi_{PR}}=\tfrac{1}{2}$, we have $D_{T}=15$ and $W_{T}=16$, so the uniform haircut factor is $\alpha=\theta_{\pi_{PR}}D_{T}/W_{T}=15/32$ and each winning account’s equity is scaled by $(1-\alpha)=17/32$, giving $e_{T}=(170/32,85/32,17/32,-3,-12)$. In particular, the ranking of positive-equity accounts is preserved under pro-rata, since all winners are multiplied by the same constant factor, illustrating its rank-preservation property.
@@ 2.5 ADL Trilemma
To define the trilemma, we have to first formally describe what constitutes exchange revenue. Recall from §2.3 that, given parameters $\alpha,\beta,\eta\in[0,1]$, the insurance fund $IF_{t}$ evolves by
$IF_{t+1}=IF_{t}+\alpha\sum_{j\in T_{t}}\tau_{t}(\Delta q_{j})+\eta\,p_{t}V_{t}+\beta\sum_{i=1}^{n}|\gamma_{t}|\,p_{t}|q_{i}|-\min\{IF_{t},D_{t}\},$
@@
Most exchanges fully collect liquidation and trading fees and potentially collect fees on funding rates. [^10] In particular, we note that HLP-like systems where fees are charged on quoting offsetting positions can be viewed as a form of exchange revenue [HyperliquidHLPVaults]. Moreover, a number of perpetuals AMM DEXs collect protocol revenue from funding flows [GMXDocs, PerpV2Docs].
@@
    Fairness: Moral hazard is bounded, i.e.,, $\omega^{\pi}_{T}/D^{\pi}_{T}=\Theta(1)$ and $\omega^{\pi}_{T}/\Delta^{\pi}_{T}=\Theta(1)$.
@@
A formal statement with precise definitions and a complete proof of the proposition appears in Appendix I.
@@
  • Solvency Ratios (§5, Appendix B). We analyze the ratio $\omega^{\pi}_{T}/D^{\pi}_{T}$ under heavy-tailed assumptions. Theorem B.3 proves that for any budget-balanced static ADL with severity $\theta_{n}$, the ratio scales as $b_{n}/(\theta_{n}n)$. Consequently, enforcing rare breaches (large $\theta_{n}$) drives the ratio to zero, violating (F).
  • Fairness Models (§6–§7, Appendices CE.2). We formalize fairness in two regimes. First, for a single round (§6), we show that capped pro-rata is the unique policy satisfying Sybil resistance and monotonicity. Second, under external shocks (§7), we prove that leverage-weighted rules (RAP) minimize future shortfall but sacrifice the most systemically important winners, highlighting the tension between short-term robustness and long-term participation.
  • Revenue and Price of Anarchy (§8, Appendix F.3). We model the revenue-solvency tradeoff as a Stackelberg game between the exchange and traders. We show that static severities yield an unbounded Price of Anarchy ($J_{n}(\pi)\sim\theta_{n}n/b_{n}$) compared to dynamic policies. This confirms that achieving (R) and (S) simultaneously requires dynamic intervention, as static rules cannot balance the trilemma asymptotically.
@@ 3.1.1 Traditional Risk Metrics
We briefly review standard metrics used in finance and regulation [Boyd2017MultiPeriodTrading, BCBS2019FRTB]. For a loss $X\geq 0$ and confidence level $\alpha\in(0,1)$:
@@ 3.1.2 ADL-Specific Efficiency Metrics
To capture the trade-off between solvency and trader welfare, we define two ratios that normalize winner survival by the scale of the default:
@@
    $\mathsf{PTSR}_{T}(\pi)=\mathop{\bf E{}}\left[\frac{\omega^{\pi}_{T}}{D^{\pi}_{T}}\right].$
    This measures the survival of the top winner relative to the total socialized loss. A low PTSR indicates that the most profitable trader is being disproportionately cannibalized to cover deficits.
@@
    $\mathsf{PMR}_{T}(\pi)=\mathop{\bf E{}}\left[\frac{\omega^{\pi}_{T}}{\Delta^{\pi}_{T}}\right].$
    This compares the top winner’s equity to the largest single loser’s shortfall, capturing the concentration of risk on both sides of the trade.
These ratios mirror the VaR/ES distinction: PTSR captures aggregate efficiency, while PMR captures tail concentration. We formalize this connection and derive exact relationships in Appendix B.
@@ 3.2 Fairness and Distributional Comparisons
While scalar metrics like PTSR provide summary statistics, they cannot fully capture the fairness of a policy across the entire population of winners. Two policies might achieve similar solvency but distribute the pain very differently—one by wiping out a few large winners (Queue), another by shaving everyone slightly (Pro-Rata). To rank policies robustly, we use tools from majorization theory [MarshallOlkinArnold2011].
@@
We compare haircut vectors $h\in{\mbox{\bf R}}^{n}$ using submajorization ($\prec_{w}$). We say a policy $\pi_{A}$ is more fair (or less concentrated) than $\pi_{B}$ if its haircut vector is submajorized by $\pi_{B}$’s, i.e., $h(\pi_{A})\prec_{w}h(\pi_{B})$. This implies that for any convex cost function $\phi$ (representing trader disutility), the aggregate distress is lower under $\pi_{A}$:
@@
This gives us a powerful, parameter-free way to claim that Pro-Rata is “fairer” than Queue: it minimizes the collective pain for all convex risk attitudes. We provide a detailed treatment of these orderings and their application to ADL in Appendix E.2.
@@
In our worst-case analysis, we often consider comonotonic couplings, where winners’ profits and losers’ deficits are perfectly correlated (move in lockstep). This represents the most dangerous regime for ADL, as large deficits coincide with large winner profits, testing the policy’s ability to extract liquidity without destroying the best traders. Our negative results in §5 exploit this structure to derive tight bounds on the Trilemma. See Appendix E.2 for further details on comonotonic risk bounds.

### v3

@@ 2.4 Autodeleveraging
    For $i\in[n]$, define $A_{i}=(B-\sum_{j=1}^{i}w_{T}(\mathfrak{p}_{\sigma(j)}))_{+}$, where $w_{T}(\mathfrak{p}_{\sigma(j)})$ is the haircutable endowment of the $j$th largest position under the ranking $\sigma$
@@
    Define $h\in[0,1]^{n}$ (or equivalently, $x\in[0,w_{T}]^{n}$) as
    $h_{\pi_{Q},\sigma(i)}=\begin{cases}1&\text{if }A_{i-1}-A_{i}=w_{T}(\mathfrak{p}_{\sigma(i)}),\;A_{i}>0\\ \frac{A_{i-1}}{w_{T}(\mathfrak{p}_{\sigma}(i))}&\text{if }A_{i}=0,A_{i-1}>0\\ 0&\text{otherwise}\end{cases}$ (23)
Note this definition of $h_{\pi_{Q}}$ ensures that $\sum_{i=1}^{n}h_{\pi_{Q},i}w_{T}(\mathfrak{p}_{i})=B$, satisfying (19). On the other hand, since we only subtract from positions with positive endowment (e.g. $j$ such that $w_{T}(\mathfrak{p}_{\sigma(j)})>0$), we satisfy (20) by construction.
@@
This haircut strategy is even simpler than the Queue and has some benefits in terms of fairness, as we will show in §6. The Pro-Rata strategy, simply put, haircuts each winning user with positive endowment by the same multiplicative fraction. The fraction is subject to budget constraints, as we describe below.
Important distinction: The pro-rata rule defined here is wealth-pro-rata (losses proportional to endowment/equity). Some exchanges (e.g., Drift) implement contracts-pro-rata (losses proportional to position size in contracts). These are different mechanisms with different fairness meanings and should not be conflated. The paper’s fairness analysis uses wealth-pro-rata as the benchmark.
Given a valid feasible shortfall $\theta_{\pi}D_{T}(\mathcal{P}_{n})\leq U_{T}(\mathcal{P}_{n})$, the wealth-pro-rata haircut policy is defined as:
$h_{\pi_{PR}}(\mathfrak{p})=\frac{\theta_{\pi}D_{T}(\mathcal{P}_{n})}{U_{T}(\mathcal{P}_{n})}\quad\text{for }w_{T}(\mathfrak{p})>0,$ (24)
or equivalently, $x_{\pi_{PR},i}=h_{\pi_{PR}}w_{i,T}$ where $h_{\pi_{PR}}=\theta_{\pi}D_{T}/U_{T}$ is the uniform haircut fraction. This ranking simply says that the haircutable endowment is socialized proportional to its size. Some exchanges slightly modify this formula to include leverage (e.g. Binance’s ADL Priority Index includes leverage and unrealized PNL [21]; Aevo documentation discusses leverage-weighted priority [4]), which penalizes higher leverage positions much like the PNL-leverage ranking:
$h_{\pi_{LPR}}(\mathfrak{p})=\left(\frac{\ell w_{T}(\mathfrak{p})}{\sum_{\mathfrak{p}\in\mathcal{P}_{n}}\ell w_{T}(\mathfrak{p})}\right)\frac{\theta_{\pi}D_{T}(\mathcal{P}_{n})}{\sum_{\mathfrak{p}\in\mathcal{P}_{n}}\ell w_{T}(\mathfrak{p})}$ (25)
@@
Figure 2: ADL severity example comparing queue and pro-rata coloring. Purple shading equals the negative equity mass (deficit) while blue shading shows the haircut mass allocated to winning traders. The queue panel’s dashed blue block at rank 2 highlights residual equity when the queue method allows partial closures; exchanges that close winners fully (e.g. Hyperliquid) would shave this bar completely. Haircut mass matches deficit mass in each panel, illustrating severity $\theta=0.50$.
@@
This corresponds to an exchange guaranteeing that a user will never lose more than $\overline{h}_{i}$% of their haircutable endowment in a single ADL round. Moreover, an exchange might also offer an absolute guarantee to users: if your equity is positive and sufficiently large, you will never have your equity cut to below some threshold $\underline{e}_{i}$. This can be represented by the constraint
$e^{\prime}_{i,T}=c_{i,T}+\mathrm{PNL}_{i,T}-h_{\pi,i}w_{i,T}\geq\underline{e}_{i}$ (27)
For PNL-only haircuts ($w_{i,T}=(\mathrm{PNL}_{i,T})_{+}$), this becomes $c_{i,T}+\mathrm{PNL}_{i,T}-h_{\pi,i}(\mathrm{PNL}_{i,T})_{+}\geq\underline{e}_{i}$. Numerous exchanges offer one or both of these guarantees on a per-ADL-round basis, including Hyperliquid, BitMEX, and Aevo [69, 25, 4].
Invariance to protected cash.
A key observation is that adding a common protected cash component $c$ to all winners is an affine shift that does not affect allocation-ordering theorems. Specifically, if two policies $A$ and $B$ allocate seized amounts $x^{A},x^{B}$ from endowment vectors $w$, then the post-ADL equity difference is:
$e^{\prime A}-e^{\prime B}=(c+\mathrm{PNL}-x^{A})-(c+\mathrm{PNL}-x^{B})=x^{B}-x^{A},$
which depends only on the endowment allocation, not on $c$. This means any theorem whose proof uses only (i) budget balance in the endowment ($\sum_{i}x_{i}=H$), and (ii) ordering/convexity/majorization properties of the survivor vector wxw-x, is unchanged after substituting the endowment vector from $(e_{T})_{+}$ to $(\mathrm{PNL}_{T})_{+}$. The cash component only affects feasibility/coverability: under PNL-only haircuts, $\mathsf{Solv}_{T}=W_{T}-D_{T}>0$ does not imply $D_{T}\leq U_{T}$.
@@
Consider five accounts with equities $e_{T}=(10,5,1,-3,-12)$, cash components $c_{T}=(2,1,0.5,0,0)$, and PNL components $\mathrm{PNL}_{T}=(8,4,0.5,-3,-12)$. Under PNL-only haircuts, the endowment vector is $w_{T}=(8,4,0.5,0,0)$ and the deficit is $D_{T}=15$ (from losers with negative equity). Let the queue policy choose severity $\theta_{\pi_{Q}}=\tfrac{1}{2}$, so the haircut budget is $B=\theta_{\pi_{Q}}D_{T}=7.5$. With queue order $\sigma=(2,1,3,5,4)$ induced by the PNL-leverage scores, the construction in (23) yields $A=(3.5,0,0,0,0)$ and haircuts $h=(0.4375,1,0,0,0)$ (seizing $x=(3.5,4,0,0,0)$). Post-ADL equities are $e^{\prime}_{T}=(2+8-3.5,1+4-4,0.5+0.5-0,-3,-12)=(6.5,1,1,-3,-12)$. Under the pro-rata rule (24) with the same severity $\theta_{\pi_{PR}}=\tfrac{1}{2}$, we have $D_{T}=15$ and $U_{T}=12.5$, so the uniform haircut factor is $h_{\pi_{PR}}=\theta_{\pi_{PR}}D_{T}/U_{T}=15/25=0.6$ and seized amounts are $x=(4.8,2.4,0.3,0,0)$, giving post-ADL equities $e^{\prime}_{T}=(2+8-4.8,1+4-2.4,0.5+0.5-0.3,-3,-12)=(5.2,2.6,0.7,-3,-12)$. In particular, the ranking of positive-equity accounts is preserved under pro-rata, since all winners’ endowments are multiplied by the same constant factor (10.6)=0.4(1-0.6)=0.4, illustrating its rank-preservation property.
@@ 2.5 ADL Trilemma
To define the trilemma, we have to first formally describe what constitutes exchange revenue. Recall from §2.3 that, given parameters $\alpha,\beta,\eta\in[0,1]$, the insurance fund $\mathsf{IF}_{t}$ evolves by
$\mathsf{IF}_{t+1}=\mathsf{IF}_{t}+\alpha\sum_{j\in T_{t}}\tau_{t}(\Delta q_{j})+\eta\,p_{t}V_{t}+\beta\sum_{i=1}^{n}|\gamma_{t}|\,p_{t}|q_{i}|-\min\{\mathsf{IF}_{t},D_{t}\},$
@@
Most exchanges fully collect liquidation and trading fees and potentially collect fees on funding rates. [^10] In particular, we note that HLP-like systems where fees are charged on quoting offsetting positions can be viewed as a form of exchange revenue [71]. Moreover, a number of perpetuals AMM DEXs collect protocol revenue from funding flows [60, 94].
@@
    Fairness: Moral hazard is bounded, i.e.,, $\upsilon^{\pi}_{T}/D^{\pi}_{T}=\Theta(1)$ and $\upsilon^{\pi}_{T}/\Delta^{\pi}_{T}=\Theta(1)$, where $\upsilon^{\pi}_{T}$ is the maximum post-ADL haircutable endowment.
@@
A formal statement with precise definitions and a complete proof of the proposition appears in Appendix J.
Scope note.
The ADL trilemma is a statement about an abstract policy model (static ADL families under Assumptions J.1–J.3). It does not rely on any empirical “two-party zero-sum fill” identity or on observing a complete execution/settlement ledger. Empirical sections map public observations onto model objects; that mapping is separate from the theorem.
Regime boundary.
The trilemma is conditional on Assumption J.3 (the structural deficit regime $\mu_{-}>\mu_{\Phi}$). It is silent when $\mu_{-}\leq\mu_{\Phi}$: in low-leverage or light-tailed regimes where expected shortfall rates are smaller than the maximum sustainable fee diversion rate, insurance can cover deficits without ADL binding, and all three goals (solvency, fairness, revenue) may be simultaneously achievable. Here, $\mu_{\Phi}$ can be interpreted as the largest fee diversion rate that remains compatible with non-declining long-run venue value (e.g., via an LTV sensitivity constraint). This is an interpretation of the regime boundary, not a change to the formal assumption used in the proof. The paper focuses on the high-leverage structural deficit regime ($\mu_{-}>\mu_{\Phi}$) because that is where major perp venues actually operate and where ADL events are empirically concentrated.
@@
  • Solvency Ratios (§5, Appendix C). We analyze the ratio $\upsilon^{\pi}_{T}/D^{\pi}_{T}$ under heavy-tailed assumptions, where $\upsilon^{\pi}_{T}$ is the maximum post-ADL endowment. Theorem C.3 proves that for any budget-balanced static ADL with severity $\theta_{n}$, the ratio scales as $b_{n}/(\theta_{n}n)$. Consequently, enforcing rare breaches (large $\theta_{n}$) drives the ratio to zero, violating (F).
  • Fairness Models (§6–§7, Appendices DF.2). We formalize fairness in two regimes. First, for a single round (§6), we show that capped pro-rata is the unique policy satisfying Sybil resistance and monotonicity. Second, under external shocks (§7), we prove that leverage-weighted rules (RAP) minimize future shortfall but sacrifice the most systemically important winners, highlighting the tension between short-term robustness and long-term participation.
  • Revenue and Price of Anarchy (§8, Appendix G.3). We model the revenue-solvency tradeoff as a Stackelberg game between the exchange and traders. We show that static severities yield an unbounded Price of Anarchy ($J_{n}(\pi)\sim\theta_{n}n/b_{n}$) compared to dynamic policies. This confirms that achieving (R) and (S) simultaneously requires dynamic intervention, as static rules cannot balance the trilemma asymptotically.
@@ 3.1.1 Traditional Risk Metrics
We briefly review standard metrics used in finance and regulation [28, 19]. For a loss $X\geq 0$ and confidence level $\alpha\in(0,1)$:
@@ 3.1.2 ADL-Specific Efficiency Metrics
To capture the trade-off between solvency and trader welfare, we define two ratios that normalize winner endowment survival by the scale of the default:
@@
    $\mathsf{PTSR}_{T}(\pi)=\mathop{\bf E{}}\left[\frac{\upsilon^{\pi}_{T}}{D^{\pi}_{T}}\right],$
    where $\upsilon^{\pi}_{T}=\max_{i}(w_{i,T}-x_{\pi,i})_{+}$ is the maximum post-ADL haircutable endowment. This measures the survival of the top winner’s endowment (profit capacity) relative to the total socialized loss. A low PTSR indicates that the most profitable trader is being disproportionately cannibalized to cover deficits.
@@
    $\mathsf{PMR}_{T}(\pi)=\mathop{\bf E{}}\left[\frac{\upsilon^{\pi}_{T}}{\Delta^{\pi}_{T}}\right].$
    This compares the top winner’s post-ADL endowment to the largest single loser’s shortfall, capturing the concentration of risk on both sides of the trade.
These ratios mirror the VaR/ES distinction: PTSR captures aggregate efficiency, while PMR captures tail concentration. Under PNL-only haircuts ($w=(\mathrm{PNL})_{+}$), these metrics measure profit capacity survival, not total equity survival. We formalize this connection and derive exact relationships in Appendix C.
@@ 3.2 Fairness and Distributional Comparisons
While scalar metrics like PTSR provide summary statistics, they cannot fully capture the fairness of a policy across the entire population of winners. Two policies might achieve similar solvency but distribute the pain very differently — one by wiping out a few large winners (Queue), another by shaving everyone slightly (Pro-Rata). To rank policies robustly, we use tools from majorization theory [84].
@@
We compare haircut vectors $h\in{\mbox{\bf R}}^{n}$ using submajorization ($\prec_{w}$). We say a policy $\pi_{A}$ is more fair (or less concentrated) than $\pi_{B}$ if its haircut vector is submajorized by $\pi_{B}$’s, i.e., $h(\pi_{A})\prec_{w}h(\pi_{B})$. A function $\phi:{\mbox{\bf R}}^{n}\to{\mbox{\bf R}}$ is said to be Schur-convex if $x\prec_{w}y$ implies $\phi(x)\leq\phi(y)$, i.e., it increases with concentration. This implies that for any convex, symmetric cost function $\phi$ (representing trader disutility), the aggregate distress is lower under $\pi_{A}$:
@@
This gives us a powerful, parameter-free way to claim that Pro-Rata is “fairer” than Queue: it minimizes the collective pain for all convex risk attitudes. We provide a detailed treatment of these orderings and their application to ADL in Appendix F.2.
@@
In our worst-case analysis, we often consider comonotonic couplings, where winners’ profits and losers’ deficits are perfectly correlated (move in lockstep). This represents the most dangerous regime for ADL, as large deficits coincide with large winner profits, testing the policy’s ability to extract liquidity without destroying the best traders. Our negative results in §5 exploit this structure to derive tight bounds on the Trilemma. See Appendix F.2 for further details on comonotonic risk bounds.